OpenStudy (anonymous):

Which of the following is an equation of the line perpendicular to 3x + 6y = 12 and passing through (4, 0)? *help please*

OpenStudy (anonymous):

Rearrange equation into standard form by dividing through by 6 so 3x + 6y = 12 is now written 1/2x + y = 2 Now make y the subject y = 2 -1/2x A perpendicular to this line will have a gradient -2 (inverse negative of original gradient) so the line will be y = -2x + c (we now have to find the constant) Put your coordinate values into this equation. 0 = -2(4) + c 0 = -8 + c Therefore c = 8 (adding 8 to both sides) Answer y = 8 -2x or 2x +y -8 = 0

OpenStudy (anonymous):

thanks darling your AMAZING (:

OpenStudy (anonymous):

Bailey I noticed an error on my previous post. All is ok up to here... Please take care to look at the changes from -2 to 2 A perpendicular to this line will have a gradient 2 (inverse negative of original gradient) so the line will be y = 2x + c (we now have to find the constant) Put your coordinate values into this equation. 0 = 2(4) + c 0 = 8 + c Therefore c = -8 (subtracting 8 to both sides) Answer y = 2x - 8 Sorry for this silly mistake, I hope it is not to late for you.